OpenStudy (anonymous):

2 girls live 3 miles apart on the same road. If they walk towards each other at rates that differ by 1/2 mph and meet exactly 24 minutes later, how fast was each girl walking?

OpenStudy (tkhunny):

How fast is Girl #1 Walking? Is she the faster or the slower? Give her speed a name, please.

OpenStudy (anonymous):

I don't know how fast Girl#1 is walking, but she is going 1/2 mph slower than girl#2.

OpenStudy (tkhunny):

Good. Girl #2 is the faster. That is an important choice. It doesn't matter which one, but you do have to pick on in order to stay organized? You did not give the NAME of the Speed of Girl #1. It is quite laborious to call it "Speed of Girl #1". We need something more convenient. What shall it be?

OpenStudy (anonymous):

You can just use x and y

OpenStudy (tkhunny):

There's a reason why I asked for only the speed of Girl #1. We could use x and y, but this might be excessive. If the speed of Girl #1 is x mph, Then the speed of Girl #2 is (x+1/2) mph. Do you agree with this assessment?

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

How long did they walk?

OpenStudy (anonymous):

For 24 minutes

OpenStudy (tkhunny):

We have Distance = Rate * Time. We're going to want "Rate * Time". Can we do that without any trouble or do you see a small problem?

OpenStudy (anonymous):

minutes and hours

OpenStudy (tkhunny):

Oh, you're good!!! We need to convert minutes to hours, since our speeds are defined in terms of hours. 24 min = What hours?

OpenStudy (anonymous):

.4 or 2/5

OpenStudy (tkhunny):

Perfect. We are now ready to solve the problem. Distance = 3 miles It's in the problem statement. Rate = (x)+(x+(1/2)) mph Since they are walking directly towards each other, we can just add them. Time = (2/5) hour This is sort of in the problem statement. We did have to translate it a little. Solve!!

OpenStudy (anonymous):

I don't understand this

OpenStudy (tkhunny):

Do we have the formula Distance = Rate * Time? Do you understand how that works?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

The fundamental idea is that if we have any two of these items, we can solve for the third. If we have Distance and Rate, we can use division and solve for Time. If we have Distance and Time, we can use division and solve for Rate. If we have Rate and Time, we can use multiplication to solve for Distance. Still with me?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

The problem statement gave us the time (24 min = 4/5 hours) and the distance (3 miles). Use this information to calculate the total speed.

OpenStudy (anonymous):

2.4

OpenStudy (tkhunny):

Try again. Distance = Rate * Time 3 miles = Rate * (4/5 hour) Rate = (3 miles) / (4/5 hour) = (3 / (4/5)) miles/hour = (3 * 5/4) mph = 3.75 mph Agreed?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

The TOTAL speed is 3.75 mph One is walking at x mph The other is walking at x + ½ mph Using this 'x' that we have defined, what is the total speed?

OpenStudy (anonymous):

so the slower person would be going 3.5 mph, and the faster would be going 4 mph.

OpenStudy (anonymous):

thank you

OpenStudy (tkhunny):

No, that's not it. The total speed would then by 7.5 mph. That's no good. We alredy know the total speed is 3.75 mph. Thus, x + (x + ½) = 3.75 You must solve this for 'x'!

OpenStudy (anonymous):

all I needed to know was how fast each girl was going. and it was right.

OpenStudy (tkhunny):

... not if it accepted 3.5 and 4.

OpenStudy (tkhunny):

* Correction. I'm not sure how it changed from 2/5 to 4/5. Missed a reasonableness check, somewhere. 3/(2/5) = 15/2 = 7.5 and we DO have 3.5 and 4.0. Not sure how the OP got there.